OpenStudy (anonymous):

Your parents have finally decided to install an in-ground pool! They come to you for help deciding between a circular pool with a 30-foot diameter and a 16 ft x 40 ft rectangular pool. The circular pool would have a 4-foot walkway and the rectangular pool would have a 3-foot walkway around the pool. The backyard measures 75 feet by 50 feet.

OpenStudy (anonymous):

@TheBleep

OpenStudy (anonymous):

Calculations for the area of the circular pool swimming area. this is the first question

OpenStudy (anonymous):

i can find tht hole answer for tht question tht you posted

OpenStudy (anonymous):

well answer it xD

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so what do I calculate first

OpenStudy (anonymous):

it is a rectangle pool 16 x 40 with a 3 foot walkway would mean that the area the pool would cover would be (3ft+16ft+3ft) * (3ft+40ft+3ft) = 22ft * 46ft = 1012 ft^2 well theres tht )

OpenStudy (anonymous):

Calculations for the area of the circular pool with the 4-foot walkway. this is question two

OpenStudy (anonymous):

grassy area is 75ft * 50ft = 3750ft^2 the area covered by the circular pool would have a diameter of 4ft+30ft+4ft=38ft the area of a circle is πr^2 : Your radius is 38ft/2 = 19ft so the area of the circular pool is π * (19ft)^2 =1134.115ft^2 the grassy area left with the circular pool is 3750ft^2-1134.115ft^2=2615.885ft^2 the grassy area left with the rectangular pool is 3750ft^2 -1012ft^2=2738ft^2

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

Calculations for the area of the rectangular pool swimming area.

OpenStudy (anonymous):

is this the third question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

:D

OpenStudy (anonymous):

how do we calculate this question

OpenStudy (anonymous):

hold on im thinking ' :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

idk about this one.. just do your best guess but the other two are good right?

OpenStudy (anonymous):

yes I have 4 more to go

OpenStudy (anonymous):

k can you give me a medal.. and also post them so i can get my problem solving better :)

OpenStudy (anonymous):

Calculations for the area of the rectangular pool with the 3-foot walkway.

OpenStudy (anonymous):

hey i g2g :/

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

just guees or ask for help can you give me a medal ?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i gave you one too :)

OpenStudy (anonymous):

yay !

OpenStudy (anonymous):

yup we both get something :)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

do you know how to give medals?